OpenStudy (anonymous):

simplify into radical form √-48

OpenStudy (anonymous):

youre hot

OpenStudy (anonymous):

\[\sqrt{-48}=\sqrt{16\times 3\times -1}=4\sqrt{3}i\]

OpenStudy (anonymous):

You have a negative number under the radical therefore your answer is going to be complex. \[\sqrt{-48} = \[i \sqrt{48} = i \sqrt{4^{2}*3} \] Find the biggest perfect square that is a factor or 48? which is 16 = 4^2

OpenStudy (anonymous):

i guess is skipped a step \[\sqrt{16\times 3\times -1}=\sqrt{16}\times \sqrt{3}\times \sqrt{-1}=4\sqrt{3}i\]

OpenStudy (anonymous):

Thank you @philo1234 @satellite73 and @bigrobot69 lol